OpenStudy (anonymous):

A gardener wants to fence in her plot in two equal rectangular sections. If she has 15 yards of fence and the area of the entire plot is 6 yd 2, find the possible dimensions in yards of the garden

OpenStudy (anonymous):

3y + 2x = 15 y*x = 6 3y +12/y = 15 y^2 - 5y + 4 = 0 (y-1)(y-4) = 0 y = {1, 4} if y = 1, x = 6 if y = 4, x = 1.5

OpenStudy (anonymous):

plz help guys i have a ton of these

OpenStudy (anonymous):

are the last the dimensions?